Harlequin Co. adopted the dollar-value LIFO retail method at the beginning of 2021 (its base year). Its beginning inventory for 2021 was $38,500 at cost and $74,500 at retail prices. At the end of 2021, it computed its estimated ending inventory at retail to be $115,000. Assuming its cost-to-retail percentage for 2021 transactions was 40%, and that the retail price index at the end of 2021 was 1.3, what is the inventory balance that Harlequin Co. would report in its 12/31/2021 balance sheet?
Answer #1
COST TO RETAIL PERCENTAGE 40%
PRICE INDEX (1.3*40%) 0.52
OPENING INVENTORY VALUE $    38,500.00
CLOSING INVENTORY VALUE $    46,000.00
ENDING VALUE AS PER INFLATION
= 46000 / 0.52 $    88,461.54
= 88461.54 - 46000 $    49,961.54
GROSS UP VALUE $    25,980.00
=49961.54 * 0.52
dollar value lifo method
OPENING INVENTORY VALUE $    38,500.00
GROSS UP VALUE $    25,980.00
dollar value lifo method $    64,480.00